# Non-Markovian Dynamics in Fiber Delay-Line Buffers

### v2

@@ Abstract
We study the non-Markovian effect on a two-photon polarization entangled state in which one photon from the pair is stored in a fiber delay-line buffer. We propose a model of a photonic qubit coupled to fiber birefringence and a fiber reservoir representing the environment. We analytically derive a non-Markovian probability function for the buffered photon and its paired photon. To verify the probability function, we perform the full quantum state tomography of the photon pairs. The probability function fits well with the experimental data and the physical values. Our result indicates that our quantum system is operating slightly above the threshold for a non-Markovian transition. We observe a unique polarization dynamic of the buffered photon. We further exploit the measures of quantum mutual information for studying the quantumness of the photon pairs. We find that Werner’s well-known separability criterion occurs at the buffer time of about 0.9 ms. Our result implies that the quantum discord can surpass Werner’s criterion, and hence the quantum bi-partite correlation can exist for a buffer time greater than 0.9 ms.
@@ I Introduction
Photons that traverse in optical fiber with little attenuation are ideal carriers for quantum networking. The photonic qubits with entanglement in polarization, time-bin, or other degrees of freedom are the resources for establishing quantum communication and distributed quantum computing. Quantum buffers made of optical fiber and free-space optics can provide the entire telecom C-band optical bandwidth and storage time up to $\mu\rm{sec}$ without significant deterioration of storage efficiency kflee24 ; pittman02 ; pittman23 ; XLi_OL_2005 ; Kwiat_SPIE_2021 ; Kwait_optica_2017 ; Wang_2022 ; Takesure_ncomm_2013 ; Park_OE_2019 ; Clemmen_OL_2018 ; Mendoza_Opa_2016 . Short buffer times (few nanoseconds) can provide active temporal source multiplexing for high-efficiency single photon generation Park_OE_2019 ; Kaneda_2015 ; Xiong_NC2016 . Long storage times play an important role in optical networks for packet synchronization, label processing, and contention management. Quantum buffers based on fiber delay lines are traveling buffersZhong_2001 ; Kaneda_2015 ; Xiong_NC2016 and fiber loop buffers Weber_1996 ; kflee24 . The traveling buffers, which can provide qubit storage time up to $\mu\rm{sec}$, are relatively unexplored. In general, there is a strong coupling between qubits, fiber birefringence, and a fiber reservoir, so we need to develop a non-Markovian model to study the decoherence dynamics of qubits in a fiber-based quantum buffer.
@@
Recent studies Giovannetti_PRL_22 ; Giovannetti_JPA_05 ; Helsen_2023 indicate that Markovian and non-Markovian processes can be controlled for improving the quantum capacity in a high-loss communication channel and modeling quantum networks. Interestingly, one can protect quantum entanglement by using the memory effect of the classical noises in the non-Markovian regimeWenjie_2023 . The fiber-based quantum buffer is an open quantum system because the polarization degree of freedom of photon is coupled with the frequency degree of freedom of the fiber birefringence which is coupled with the fiber reservoir. The density matrix of a qubit can evolve into Markovian (non-Markovian) dynamics depending on the weak (strong) coupling with the medium, respectively. The medium here is a single-mode optical fiber birefringence and a fiber reservoir. In a weak coupling process, the medium induces memoryless dynamics on the qubit so that its density matrix $\rm{\rho(t)}$ at time t is determined by $\rm{\rho(t-dt)}$ with the infinitesimal $dt$. In other words, the information of a qubit should not be memory-related with the previous qubit. The Markovian dynamic of a qubit is ideal for long-distance communication networks. The non-Markovian process, in contrast, keeps the memory of the coherence dynamics of the qubit such that its density matrix $\rm{\rho(t)}$ at time t is determined by its past $\rm{\rho(t^{\prime})}$ where $t^{\prime}$ is the start of the strong coupling process. It is essential to minimize the non-Markovian dynamic of a qubit stored in a fiber-based buffer by exploring its coupling with fiber birefringence and fiber reservoir. A comprehensive discussion of concepts, measures, and witness of non-Markovian effect for an open quantum system can be found in refPlenio_2014 ; Piilo_16 . A trace distance of two quantum states Piilo_09 ; Piilo_10 is usually used to detect the crossover between the Markovian and non-Markovian processes such as the coupling between the photon polarization and its optical frequency degree of freedom in a simple titled Fabry-Perot cavity platform Guo_nphy_2011 . The non-Markovian effect can revive entanglement after observation of entanglement sudden death Haase_2018 ; Wang_2018 .
@@
An exactly solvable master equation of an open quantum system, such as a two-level atom interacting resonantly in a reservoir formed by the Lorentzian quantized modes of a high Q cavity, is often used as an example for illustrating the non-Markovian dynamics of a qubit Breuer_oxford_2002 ; Petruccione_PRA_2006 ; Bellomo_PRL_2007 ; Chen_PRA_14 ; Wang_scr_17 . Furthermore, the non-Markovian dynamics have been studied by using quantum discord Wang_PRA_2010 ; Fanchini_PRA_2010 , which is another concept of quantum correlation introduced by quantum mutual information theory. Some believe that quantum discord is more suitable for quantifying the quantum correlation of a bipartite system in non-Markovian environments. We recently used it to explore the quantum coherence of a two-qubit entangled state generated in green fluorescent protein Kflee_nc_2017 ; Shi_sr_2016 at room temperature. One can experimentally obtain quantum discord by calculating the total correlation and classical correlation based on the density matrix elements of a quantum state.
@@
In this paper, one photon from the polarization entangled photon pair is stored in the buffer. We use a model to study the non-Markovianity of a photonic qubit coupled to the fiber birefringence of a single-mode optical fiber in a fiber reservoir. We analytically derive the probability function for the buffered polarization qubit and its paired photon. Quantum state tomography (QST) is then performed on the entangled photon pairs. We obtain the probability function from the reconstructed density matrix. Our model agrees with the experimental data. The buffered photon exhibits the non-Markovianity because of the polarization modal dispersion (PMD) and the loss in the fiber reservoir. In the experiment, we also prove the existence of the Markovian process in the scenario where we replace the buffer with a free-space attenuator.
@@ II Results
Figure 1: A simple model consists of a photonic qubit, fiber birefringence in optical fiber, and a fiber reservoir representing the environment decoherence. The $\kappa$ is the coupling constant between the qubit and fiber birefringence ($\Delta n$). The $\kappa(\rm{sec}^{-1})$ is inversely proportional to PMD ($\sim\frac{1}{\sqrt{\langle\tau^{2}\rangle}}$), where $\langle\tau^{2}\rangle$ is differential group delay (DGD). The fiber birefringence is then coupled with a fiber reservoir with the coupling constant $\gamma_{\circ}(\rm{sec}^{-1})$ which is responsible for the loss of qubit information such as decoherence.
@@
Here, we consider a qubit-fiber-reservoir model consisting of a photonic qubit, a single-mode optical fiber, and a fiber reservoir representing the environment as shown in Fig.1. (see Methods for the atom-cavity-reservoir model). When a photon propagates in a long optical fiber, its polarization state and transmission rate will be degraded by the fiber birefringence and a fiber reservoir. The fiber reservoir is the medium where a polarization qubit interacts/exchanges energy with the environment. The polarization of the photon with certain optical frequency bandwidth is coupled with the fiber birefringence which is coupled with the fiber reservoir. In other words, the PMD effect is coupled to the environment via the fiber reservoir. The photon polarization state will experience the dephasing caused by the PMD and the absorption/scattering process in the fiber. The absorption and re-emission of a photon state in the fiber will decohere a quantum state. A photonic qubit with polarization state H or V is coupled with the fiber birefringence by a coupling constant $\kappa\propto\frac{1}{\rm{PMD}}$. A photonic qubit in a single optical frequency mode will couple weakly with the fiber birefringence or less dispersion (large $\kappa$), and hence experiencing less depolarization effect. If we increase the frequency bandwidth of the photonic qubit in a steep dispersion region, the $\kappa$ becomes small. The qubit will start to experience the PDM effect, which is given by the square root of time variance $\langle\tau^{2}\rangle$ of differential group delay (DGD) Kogelnik_PMD_2002 ; Wai_jlt_1996 ; Brodsky_jlt_2006 or $\sqrt{\langle\tau^{2}\rangle}$. Therefore, the strength of the PMD effect is inversely proportional to the $\kappa$. In other words, one can enhance the PMD effect on the photonic qubit by increasing their differential group delays such as inserting a few polarization-maintaining fiber jumpers in their channels. The same photon with optical frequency $\omega$ can be spontaneously absorbed and scattered at a rate of $\gamma_{\circ}$ due to the coupling between the fiber birefringence and the reservoir. The $\frac{\gamma_{\circ}}{c}\propto\mu$ is the loss parameter per unit length of the photon due to photon absorption/scattering in the fiber reservoir. (see Methods for the probability function of the photon state and how the relative strength of the $\gamma_{\circ}$ and $\kappa$ can determine the photon state undergoing Markovian and non-Markovian processes). Two factors contribute to the loss parameter. The first is the propagation loss of the photonic qubit in the fiber. The second is the loss of information on the photonic qubit which is its polarization property. The environment depolarization process can degrade the polarization extinction ratio of the qubit. The propagation loss of the photonic qubit can cause an increase in the accidental coincidence and reduce the fidelity of the quantum state. In this paper, we subtract the accidental coincidence of the signal and idler so that we can isolate the contribution of the propagation loss to the reconstructed density matrix and study the polarization property of the buffered qubit which is the idler photon.
@@
Figure 2: a) H(V) analog to the excited state (ground state) of a two-level atomic system. The geometry picture for the change of polarization state (i) $\rm{H}\rightarrow\rm{H}^{f}$ and (ii) $\rm{V}\rightarrow\rm{V}^{f}$. Small arrows along the $\rm{H}^{f}$ and $\rm{V}^{f}$ show the contribution from the H and V. The phase fluctuation $\Delta\phi$ due to PMD degrades the polarization extinction ratio of the idler. b) Experiment setup including the buffer made of two low-loss switches with a short path and a long path. We use different fiber lengths in the long path to change the buffer time.
@@
In our two-qubit system, the photon pair is initially prepared in the entangled state $|\psi_{\circ}\rangle=\frac{1}{\sqrt{2}}[|H_{s}H_{i}\rangle+|V_{s}V_{i}\rangle]$. One photon from the pair is sent to a buffer with the fiber length of L. Let’s say the photon of interest is the idler which is experiencing the PMD and the loss in the buffer. The initial state of the idler (signal) is $|H_{i}\rangle+|V_{i}\rangle$ ($|H_{s}\rangle+|V_{s}\rangle$), respectively. Due to depolarizing caused by the PMD and fiber reservoir, the final components $H^{f}_{i}$ and $V^{f}_{i}$ can be written in terms of $H_{i}$ and $V_{i}$ as
@@
where $\Delta\phi_{h(v)}$ is the amount of phase change caused by the PMD effect. The phase change $\Delta\phi_{h(v)}=\Delta\omega\,\sqrt{\langle\tau^{2}\rangle}$ is related to the square root of time variance $\langle\tau^{2}\rangle$ of differential group delay (DGD) Kogelnik_PMD_2002 ; Wai_jlt_1996 ; Brodsky_jlt_2006 or $\sqrt{\langle\tau^{2}\rangle}\propto L^{1/2}$, where L is the length of the fiber buffer. The $h$ ($v$) sign denotes the phase change associated with the polarization state $H_{i}$ ($V_{i}$). When the $\Delta\phi_{h}$ and $\Delta\phi_{v}$ are zero, the $H^{f}_{i}\rightarrow H_{i}$ and $V^{f}_{i}\rightarrow V_{i}$. The polarization states of the idler do not change because of the absence of the PMD effect. When the $\Delta\phi_{h}$ and $\Delta\phi_{v}$ are +90(90)9090+90(-90)+ 90 ( - 90 ), $H^{f}_{i}\rightarrow V_{i}(-V_{i})$ and $V^{f}_{i}\rightarrow-H_{i}(+H_{i})$, respectively. The polarization states of the idler are completely flipped. The geometrical meaning for these notations is shown in Fig.2(a). The phase fluctuation of $\rm{cos(\Delta\phi)}$ and $\rm{sin(\Delta\phi)}$ degrade the polarization extinction ratio of the idler. One can interpret that the final idler polarization state $H^{f}_{i}$ is related to its initial states $H_{i}$ and $V_{i}$ in the past because of the non-Markovian effect caused by the PMD. We assume that the signal is propagating in a short fiber length before it is detected. We then obtain the probability function $\mathcal{P}_{f}=|\langle\psi_{\circ}|\psi^{f}\rangle|^{2}$ as,
@@
where $\mathcal{A}\mathcal{A}^{*}=e^{-2\mu L}$ is introduced due to the loss parameter in the fiber reservoir on the polarization properties experienced by the surviving idler photons. In our experiment, we diminish the contribution of the propagation loss to the probability function by subtracting the accidental coincidence, so that we can measure the probability function which is degraded by the $\mathcal{A}\mathcal{A}^{*}=e^{-2\mu L}$ due to the fiber reservoir.
We explore non-Markovian dynamics of the entangled state by simplifying Eq.(2) into two cases; the magnitude of $\Delta\phi_{h}=\Delta\phi_{v}$ and $\Delta\phi_{h}\neq\Delta\phi_{v}$. For the case $\Delta\phi_{h}=\Delta\phi_{v}=\Delta\phi$, we have the first scenario where the idler polarization states $H_{i}$ and $V_{i}$ rotate in opposite directions i.e we substitute the terms $+\Delta\phi_{h}$ and $-\Delta\phi_{v}$ or $-\Delta\phi_{h}$ and $+\Delta\phi_{v}$ in Eq.(2). The $+$ (--) value of the $\Delta\phi_{h(v)}$ corresponds to the rotation of the polarization state in the clockwise (anti-clockwise) direction. The probability function $\mathcal{P}_{f}$ is simplified as
@@
where $\Delta\phi_{1}=\Delta\omega\,\sqrt{\langle\tau_{1}^{2}\rangle}$. Eq.(3) has a similar form as Eq.(8) in Methods for a two-level atom interacting in a high Q-cavity with a memoryless reservoir. This implies that the idler is experiencing the non-Markovian process due to the coupling between the PMD effect and the loss in the fiber reservoir. Let’s consider the second scenario where the idler polarization states $H^{f}_{i}$ and $V^{f}_{i}$ rotate in the same directions i.e, we substitute the terms $+\Delta\phi_{h}$ and $+\Delta\phi_{v}$ or $-\Delta\phi_{h}$ and $-\Delta\phi_{v}$ in Eq.(2). Then the probability function $\mathcal{P}_{f}$ becomes
@@
where $\Delta\phi_{2}=\Delta\omega\,\sqrt{\langle\tau_{2}^{2}\rangle}$. The Eq.(4) has the same form as Eq.(9) in Methods for large $n$. For the case $\Delta\phi_{h}\neq\Delta\phi_{v}$, the complete probability function is shown in detail in the Method.
@@
Later, we will verify the probability functions of Eq.(3) and Eq.(4) and also Eq.(8) and Eq.(9) in the Methods by launching one photon from the entangled photon pair in a quantum buffer made of different fiber lengths as shown in Fig.2(b). We use a counter-propagating scheme (CPS) Kflee_ol_2006 ; Sua_ol_2014 to generate two-photon polarization entangled state $|\psi_{\circ}\rangle$ through a four-wave mixing process in a 300 m of dispersion-shifted fiber. The detailed experimental setup is shown in Fig.1S in the Supplementary. The signal and idler photons are separated by the DWDMs (dense wavelength-division multiplexing). The signal photon is sent to the polarization analyzer $\rm{PA_{s}}$. The idler is sent to the buffer which consists of two low-loss switches (insertion loss $\leq$ 0.01dB and a single-mode optical fiber as shown in Fig.2(b). We can select the idler to pass a long or short fiber length. When the idler is selected to pass the long fiber, then the idler is buffered with the duration of $\frac{L}{c/n_{r}}$, where $L$ is the length of the fiber and $n_{r}$ is the refractive index of the fiber. The idler is then directed to the polarization analyzer $\rm{PA_{i}}$. The first pair of the quarter-wave plate (QWP) and half-wave plate (HWP) is used to compensate for the birefringence of the fiber. The second pair of QWP and HWP is used to implement quantum state tomography. The timing of detecting the buffered photon can cause a non-Markovian effect on the photon stateHelsen_2023 ; Humprey_Nature_18 . In this buffer configuration, the qubit storing process is initialized by two units of 1$\times$2 switch. This switching process can impose loss and decoherence such as the amplitude/phase damping process. We use two units of 1$\times$2 switch instead of one unit of 2$\times$2 switch because the insertion loss of 1$\times$2 switch (2$\times$2 switch) is about $<$ 0.01 dB (1.2 dB), respectively. To avoid decoherence due to timing issues on the photon detection, we operate the switch with low speed i.e. 1.0Hz. The slow switching rate allows us to take data when the switches are both turned on. So, we can exclude the data appears in the rising and falling edges of the switches. In other words, we diminish the non-Markovian effect which can be caused by the switching process kflee24 . We use an interface program to turn on and off the two switches synchronously at the rate of 1.0 Hz. The data is taken on the duration when both switches are completely turned on/off. We then perform quantum state tomography (QST) by using two single-photon detectors (NuCrypt CPDS-4) operated at 50 MHz.

### v3

@@ Abstract
We study the non-Markovian effect on a two-photon polarization entangled state, in which one photon from the pair is stored in a fiber delay-line buffer. We propose a model of a photonic qubit coupled to fiber birefringence and a fiber reservoir representing the environment. We analytically derive a non-Markovian probability function for the buffered photon and its paired photon. To verify the probability function, we perform full quantum state tomography of the photon pairs. The probability function fits well with the experimental data and physical values. Our results indicate that our quantum system operates slightly above the threshold for a non-Markovian transition. We observe a unique polarization dynamic of the buffered photon. We further exploit measures of quantum mutual information to study the quantumness of the photon pairs. Werner’s well-known separability criterion occurs at a buffer time of about 0.9 ms. Our results imply that quantum discord can surpass Werner’s criterion, and hence, quantum bi-partite correlation can exist for buffer times greater than 0.9 ms.
@@ I Introduction
Photons that traverse in optical fiber with little attenuation are ideal carriers for quantum networking. Photonics qubits with entanglement in polarization, time-bin, and other degrees of freedom are essential resources for establishing quantum communication and distributed quantum computing. Quantum buffers made of optical fiber and free-space optics can operate over the entire telecom C-band optical bandwidth and have storage times up to microseconds without significant deterioration of storage efficiency kflee24 ; pittman02 ; pittman23 ; XLi_OL_2005 ; Kwiat_SPIE_2021 ; Kwait_optica_2017 ; Wang_2022 ; Takesure_ncomm_2013 ; Park_OE_2019 ; Clemmen_OL_2018 ; Mendoza_Opa_2016 . Short buffer times (a few nanoseconds) can provide active temporal source multiplexing for high-efficiency single-photon generation Park_OE_2019 ; Kaneda_2015 ; Xiong_NC2016 . Long storage times play an important role in optical networks for packet synchronization, label processing, and contention management. Quantum buffers based on fiber delay lines can be traveling buffers Zhong_2001 ; Kaneda_2015 ; Xiong_NC2016 or fiber loop buffers Weber_1996 ; kflee24 . The traveling buffers, which can provide qubit storage time up to microseconds, are relatively unexplored. In general, there is a strong coupling between qubits, fiber birefringence, and a fiber reservoir, so we need to develop a non-Markovian model to study the decoherence dynamics of qubits in a fiber-based quantum buffer.
@@
Recent studies Giovannetti_PRL_22 ; Giovannetti_JPA_05 ; Helsen_2023 indicate that Markovian and non-Markovian processes can be controlled to improve the quantum capacity in high-loss communication channels and model quantum networks. Interestingly, one can protect quantum entanglement by using the memory effect of the classical noises in the non-Markovian regimeWenjie_2023 . The fiber-based quantum buffer is an open quantum system because the polarization degree of freedom of the photon is coupled with the frequency degree of freedom of the fiber birefringence, which is coupled with the fiber reservoir. The density matrix of a qubit can evolve into Markovian (non-Markovian) dynamics depending on the weak (strong) coupling with the medium, respectively. The medium here is a birefringent single-mode optical fiber and a fiber reservoir. In a weak coupling process, the medium induces memoryless dynamics on the qubit, so its density matrix $\rm{\rho(t)}$ at time t is determined by $\rm{\rho(t-dt)}$ with the infinitesimal dt. In other words, the information of a qubit should not be memory-related to the previous qubit. In this work, we study the time evolution of a two-photon state in the medium. The Markovian dynamics of a qubit are ideal for long-distance communication networks. The non-Markovian process, in contrast, retains the memory of the coherence dynamics of the qubit such that its density matrix $\rm{\rho(t)}$ at time t is determined by its past $\rm{\rho(t^{\prime})}$ where $\rm{t^{\prime}}$ is the start of the strong coupling process. It is essential to minimize the non-Markovian dynamics of a qubit stored in a fiber-based buffer by exploring its coupling with fiber birefringence and the fiber reservoir. A comprehensive discussion of concepts, measures, and witnesses of the non-Markovian effect for an open quantum system can be found in references Plenio_2014 ; Piilo_16 . A trace distance of two quantum states Piilo_09 ; Piilo_10 is usually used to detect the crossover between Markovian and non-Markovian processes, such as the coupling between the photon polarization and its optical frequency degree of freedom in a simple titled Fabry-Perot cavity platform Guo_nphy_2011 . The non-Markovian effect can revive entanglement after the observation of entanglement sudden death Haase_2018 ; Wang_2018 .
@@
An exactly solvable master equation of an open quantum system, such as a two-level atom interacting resonantly in a reservoir formed by the Lorentzian quantized modes of a high Q cavity, is often used as an example to illustrate the non-Markovian dynamics of a qubit Breuer_oxford_2002 ; Petruccione_PRA_2006 ; Bellomo_PRL_2007 ; Chen_PRA_14 ; Wang_scr_17 . Furthermore, the non-Markovian dynamics have been studied using quantum discord Wang_PRA_2010 ; Fanchini_PRA_2010 , another concept of quantum correlation introduced by quantum mutual information theory. Some believe that quantum discord is more suitable for quantifying the quantum correlation of a bipartite system in non-Markovian environments. We recently used it to explore the quantum coherence of a two-qubit entangled state generated in green fluorescent protein Kflee_nc_2017 ; Shi_sr_2016 at room temperature. One can experimentally obtain quantum discord by calculating the total correlation and classical correlation based on the density matrix elements of a quantum state.
@@
In this paper, one photon from the polarization entangled photon pair is stored in the buffer. We use a model to study the non-Markovianity of the photon pair, where the buffered qubit state is coupled to the fiber birefringence of a single-mode optical fiber in a fiber reservoir. We analytically derive the probability function for the buffered polarization qubit and its paired photon. Quantum state tomography (QST) is then performed on the entangled photon pairs. We obtain the probability function from the reconstructed density matrix. Our model agrees with the experimental data. The photon pair exhibits non-Markovianity because the buffered photon experiences the polarization modal dispersion (PMD) and loss in the fiber reservoir. In the experiment, we also prove the existence of the Markovian process in the scenario where we replace the buffer with a free-space attenuator.
@@ II Results
Figure 1: A simple model consists of a photonic qubit, fiber birefringence in optical fiber, and a fiber reservoir representing the environmental decoherence. The $\kappa$ is the coupling constant between the qubit and fiber birefringence ($\Delta n$). The $\kappa(\rm{sec}^{-1})$ is inversely proportional to PMD ($\sim\frac{1}{\sqrt{\langle\tau^{2}\rangle}}$), where $\langle\tau^{2}\rangle$ is the differential group delay (DGD). The fiber birefringence is then coupled with a fiber reservoir with the coupling constant $\gamma_{\circ}(\rm{sec}^{-1})$, which is responsible for the loss of qubit information.
@@
Here, we consider a qubit-fiber-reservoir model consisting of a photonic qubit, a single-mode optical fiber, and a fiber reservoir representing the environment as shown in Fig.1. (see Methods for the atom-cavity-reservoir model). When a photon propagates in a long optical fiber, its polarization state and transmission rate will be degraded by the fiber birefringence and a fiber reservoir. The fiber reservoir is the medium where a polarization qubit interacts and exchanges energy with the environment. The polarization of the photon with a certain optical frequency bandwidth is coupled with the fiber birefringence, which is coupled with the fiber reservoir. In other words, the PMD effect is coupled to the environment via the fiber reservoir. The photon polarization state will experience depolarization caused by the PMD and the absorption/scattering process in the fiber. The absorption and re-emission of a photon state in the fiber will decohere a quantum state. A photonic qubit with polarization state H or V is coupled with the fiber birefringence by a coupling constant $\kappa\propto\frac{1}{\rm{PMD}}$. A photonic qubit in a single optical frequency mode will couple weakly with the fiber birefringence, experiencing less dispersion (large $\kappa$), and hence less depolarization effect. If we increase the frequency bandwidth of the photonic qubit in a steep dispersion region, the $\kappa$ becomes small. The qubit will start to experience the PMD effect, which is given by the square root of time variance $\sqrt{\langle\tau^{2}\rangle}$ of differential group delay (DGD) Kogelnik_PMD_2002 ; Wai_jlt_1996 ; Brodsky_jlt_2006 . Therefore, the strength of the PMD effect is inversely proportional to the $\kappa$. In other words, one can enhance the PMD effect on the photonic qubit by increasing its differential group delay, such as by inserting a few polarization-maintaining fiber jumpers in its channel. The same photon with optical frequency $\omega$ can be spontaneously absorbed and scattered at a rate of $\gamma_{\circ}$ due to the coupling between the fiber birefringence and the reservoir. The $\frac{\gamma_{\circ}}{c}\propto\mu$ is the loss parameter per unit length of the photon due to photon absorption/scattering in the fiber reservoir. (see Methods for the probability function of the photon state and how the relative strength of the $\gamma_{\circ}$ and $\kappa$ can determine the photon state undergoing Markovian and non-Markovian processes). Two factors contribute to the loss parameter. The first is the propagation loss of the photonic qubit in the fiber. The second is the loss of information on the photonic qubit, which is its polarization property. The environmental depolarization process can degrade the polarization extinction ratio of the qubit. The propagation loss of the photonic qubit can cause an increase in accidental coincidence and reduce the fidelity of the quantum state. In this paper, we subtract the accidental coincidence of the signal and idler so that we can isolate the contribution of the propagation loss to the reconstructed density matrix and study the polarization property of the buffered qubit, which is the idler photon.
@@
Figure 2: a) H(V) analog to the excited state (ground state) of a two-level atomic system. The dynamic process for the change of the idler’s polarization state (i) $\rm{H}\rightarrow\rm{H}^{f}$ with $\Delta\phi_{h}=\Delta\phi$ and (ii) $\rm{V}\rightarrow\rm{V}^{f}$ with $\Delta\phi_{v}=-\Delta\phi$ Small arrows along the $\rm{H}^{f}$ and $\rm{V}^{f}$ show the contribution from the H and V. The phase fluctuation $\Delta\phi$ due to PMD degrades the polarization extinction ratio of the idler. b) Experiment setup including the buffer made of two low-loss switches with a short path and a long path. We use different fiber lengths in the long path to change the buffer time. Polarization analyzer (PA) consists of a quarter-wave plate (QWP), half-wave plate (HWP), and polarization beam splitter (PBS).
@@
In our two-qubit system, the photon pair is initially prepared in the entangled state $|\psi_{\circ}\rangle=\frac{1}{\sqrt{2}}[|H_{s}H_{i}\rangle+|V_{s}V_{i}\rangle]$. One photon from the pair is sent to a buffer with a fiber length of L. Let’s say the photon of interest is the idler, which is experiencing both PMD and loss in the buffer. The initial state of the idler (signal) is $|H_{i}\rangle+|V_{i}\rangle$ ($|H_{s}\rangle+|V_{s}\rangle$), respectively. Due to depolarization caused by the PMD and fiber reservoir, the final components $H^{f}_{i}$ and $V^{f}_{i}$ can be written in terms of $H_{i}$ and $V_{i}$ as
@@
where $\Delta\phi_{h(v)}$ is the amount of phase change caused by the PMD effect. The phase change $\Delta\phi_{h(v)}=\Delta\omega\,\sqrt{\langle\tau^{2}\rangle}$ is related to the square root of time variance $\sqrt{\langle\tau^{2}\rangle}$ of differential group delay (DGD) Kogelnik_PMD_2002 ; Wai_jlt_1996 ; Brodsky_jlt_2006 or $\sqrt{\langle\tau^{2}\rangle}\propto L^{1/2}$, where L is the length of the fiber buffer. The $h$ ($v$) sign denotes the phase change associated with the polarization state $H_{i}$ ($V_{i}$). The dynamic process of the idler’s polarization state described by Eq.(1) is geometrically illustrated in Fig.2(a). When $\Delta\phi_{h}$ and $\Delta\phi_{v}$ are zero, $H^{f}_{i}\rightarrow H_{i}$ and $V^{f}_{i}\rightarrow V_{i}$. The polarization states of the idler do not depolarize due to the absence of the PMD effect. When $\Delta\phi_{h}$ and $\Delta\phi_{v}$ are $+90^{\circ}(-90^{\circ})$, $H^{f}_{i}\rightarrow V_{i}(-V_{i})$ and $V^{f}_{i}\rightarrow-H_{i}(+H_{i})$, respectively. The polarization states of the idler are completely flipped. When $\Delta\phi_{h}=\Delta\phi$, the final component $H^{f}$ of the idler consists of two components: the $\rm{cos(\Delta\phi)}$ of the initial component $H_{i}$ and the $\rm{sin(\Delta\phi)}$ of the initial component $V_{i}$, as shown in Fig.2(a)(i). The $H$ component of the idler’s polarization state is decreased by an amount of $\rm{cos(\Delta\phi)}$. The polarization extinction ratio of the idler photon is degraded because of the creation of $V$ component due to the PMD effect. A similar process occurs for $\Delta\phi_{v}=-\Delta\phi$; the final component $V^{f}$ of the idler consists of the $\rm{cos(\Delta\phi)}$ of the initial component $V_{i}$ and the $\rm{sin(\Delta\phi)}$ of the initial component $H_{i}$, as shown in Fig.2(a)(ii). The phase fluctuation of $\rm{cos(\Delta\phi)}$ and $\rm{sin(\Delta\phi)}$ degrades the polarization extinction ratio of the idler. One can interpret that the final idler polarization state $H^{f}_{i}$ is related to its initial states $H_{i}$ and $V_{i}$ in the past due to the non-Markovian effect caused by the PMD. We assume that the signal is propagating in a short fiber length before it is detected. We then obtain the probability function $\mathcal{P}_{f}=|\langle\psi_{\circ}|\psi^{f}\rangle|^{2}$ as,
@@
where $\mathcal{A}\mathcal{A}^{*}=e^{-2\mu L}$ is introduced due to the loss parameter in the fiber reservoir on the polarization properties experienced by the surviving idler photons. In our experiment, we diminish the contribution of the propagation loss to the probability function by subtracting out the accidental coincidence counts so that we can measure the probability function, which is degraded by $\mathcal{A}\mathcal{A}^{*}=e^{-2\mu L}$ due to the fiber reservoir.
We explore the non-Markovian dynamics of the entangled state by simplifying Eq.(2) into two cases: the magnitude of $\Delta\phi_{h}=\Delta\phi_{v}$ and $\Delta\phi_{h}\neq\Delta\phi_{v}$. For the case $\Delta\phi_{h}=\Delta\phi_{v}=\Delta\phi$, we have the first scenario where the idler polarization states $H_{i}$ and $V_{i}$ rotate in opposite directions, i.e, we substitute the terms $+\Delta\phi_{h}$ and $-\Delta\phi_{v}$ or $-\Delta\phi_{h}$ and $+\Delta\phi_{v}$ in Eq.(2). The $+$ (--) value of the $\Delta\phi_{h(v)}$ corresponds to the rotation of the polarization state in the clockwise (anti-clockwise) direction. The probability function $\mathcal{P}_{f}$ is simplified as
@@
where $\Delta\phi_{1}=\Delta\omega\,\sqrt{\langle\tau_{1}^{2}\rangle}$. Eq.(3) is similar in structure to Eq.(8) in Methods for a two-level atom interacting in a high Q-cavity with a memoryless reservoir. This implies that the idler is experiencing the non-Markovian process due to the coupling between the PMD effect and the loss in the fiber reservoir. Let’s consider the second scenario where the idler polarization states $H^{f}_{i}$ and $V^{f}_{i}$ rotate in the same directions, i.e., we substitute the terms $+\Delta\phi_{h}$ and $+\Delta\phi_{v}$ or $-\Delta\phi_{h}$ and $-\Delta\phi_{v}$ in Eq.(2). Then the probability function $\mathcal{P}_{f}$ becomes:
@@
where $\Delta\phi_{2}=\Delta\omega\,\sqrt{\langle\tau_{2}^{2}\rangle}$. Eq.(4) has a similar form to Eq.(9) in the Methods section. For the case $\Delta\phi_{h}\neq\Delta\phi_{v}$, the complete probability function is shown in detail in the Methods section.
@@
Later, we will verify the probability functions of Eq.(3) and Eq.(4) and also Eq.(8) and Eq.(9) in the Methods by launching one photon from the entangled photon pair in a quantum buffer made of different fiber lengths, as shown in Fig.2(b). We use a counter-propagating scheme (CPS) Kflee_ol_2006 ; Sua_ol_2014 to generate the two-photon polarization-entangled state $|\psi_{\circ}\rangle$ through a four-wave mixing process in 300  meters of dispersion-shifted fiber. The detailed experimental setup is shown in Fig.1S in the Supplementary. The signal and idler photons are separated by the DWDMs (dense wavelength-division multiplexers). The signal photon is sent to the polarization analyzer ($\rm{PA_{s}}$). The idler is sent to the buffer, which consists of two low-loss switches (insertion loss $\approx$ 0.01dB) and a single-mode optical fiber, as shown in Fig.2(b). We can select the idler to pass a long or short fiber length. When the idler is selected to pass the long fiber, it is buffered for the duration of $\frac{L}{c/n_{r}}$, where $L$ is the length of the fiber and $n_{r}$ is the refractive index of the fiber. The idler is then directed to the polarization analyzer ($\rm{PA_{i}}$). We use a polarized classical light source to align the optical components in the polarization analyzers. The first pair of QWP (quarter-wave plate) and HWP (half-wave plate) is used to align the light polarization (H or V) along the principal axes of the second pair of QWP and HWP, and the polarizer The second pair of QWP and HWP is used to implement quantum state tomography. The timing of detecting the buffered photon can cause a non-Markovian effect on the photon state Helsen_2023 ; Humprey_Nature_18 . In this buffer configuration, the qubit storing process is initialized by two units of 1$\times$2 switch. This switching process can impose loss and decoherence, such as the amplitude or phase damping process. We use two units of 1$\times$2 switch instead of one unit of 2$\times$2 switch because the insertion loss of 1$\times$2 switch (0.01 dB) is much lower than that of the 2$\times$2 switch (1.2 dB). To avoid decoherence due to timing issues on the photon detection, we operate the switch at low speed, i.e., 1 Hz. The slow switching rate allows us to take data when the switches are both fully turned on. Thus, we can exclude the data that appears in the rising and falling edges of the switches. In other words, we diminish the non-Markovian effect that can be caused by the switching process kflee24 . We use an interface program to turn the two switches on and off synchronously at the rate of 1 Hz. The data is taken during the duration when both switches are completely turned on or off. We then perform quantum state tomography (QST) using two single-photon detectors (NuCrypt CPDS-4) operated at 50 MHz.